# Importance sampling for Bayesian inference: polynomial-dimension dependent error bounds

### 02 Apr 2026 v1

@@ 4.1 Linear Gaussian model
$\int_{\mathcal{Y}}\frac{\mathcal{N}({\sf d}y;\mu_{y},S_{2})}{\mathcal{N}({\sf d}y;\mu_{y},S_{1})}\,{\sf d}y=\frac{|R|^{\frac{1}{2}}}{2^{\frac{d_{y}}{2}}(2\uppi)^{\frac{d_{y}}{2}}|S_{2}|}$ (30)
and then, combining (30) and (25), we readily see that
$\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=\frac{1}{\left(2^{d_{y}}(2\uppi)^{3d_{y}}\,|R|\right)^{1/2}\,|S_{2}|}=\mathrel{\mathop{\ordinarycolon}}K_{2}^{d_{x}},$ (31)
which is always finite (provided that $R\succ 0$, which implies $S_{2}\succ 0$ as well), ensuring that the moment condition (20) holds and so does the Bochner integrability condition (13).
Moreover, the quantity in (31) can be bounded in terms of the model parameters. Indeed, recalling (22), we readily see that
$S_{2}=\Sigma_{y}-\tfrac{1}{2}R=A\Sigma A^{\top}+\tfrac{1}{2}R\succ 0,$
If we additionally observe that $S_{2}-\frac{1}{2}R\succ 0$ then, by Theorem 25 in [26],
$|S_{2}|\geq\cfrac{1}{2^{d_{y}}}|R|,\quad\text{hence}\quad\frac{1}{|S_{2}|}\leq\frac{2^{d_{y}}}{|R|},$
and the expectation in (31) can be bounded as
$K_{2}^{d_{x}}\leq\frac{1}{\,\left((2^{\frac{2}{3}}\uppi)^{d_{y}}\,|R|\right)^{\frac{3}{2}}}=\mathrel{\mathop{\ordinarycolon}}\mathcal{K}_{2}<\infty,$ (32)
that does not depend on the state dimension $d_{x}$. In other words, $K_{2}^{d_{x}}$ is bounded by a polynomial of degree zero in the sense of Corollary 3.4, i.e., $n=0$, and hence the assumptions of the corollary are satisfied.
@@
We have found that, in the linear Gaussian setting, the Bochner integrability condition (31) holds with a finite constant $\mathcal{K}_{2}$ that is independent of the state dimension. Consequently, the assumptions of Corollary 3.4 are satisfied, and the corresponding error bounds for the random IS estimator hold uniformly over $d_{x}$ and decay at the rate $\mathcal{O}(N^{-1/2})$.
4.1.1 Linear Gaussian model with arbitrary $d_{x}$ and increasing $d_{y}$
Based on the inequality (32), which holds for any $d_{x}\in\mathbb{N}$, we study the behaviour of the error bound as the observation dimension $d_{y}$ increases. Assume that there is a sequence of positive definite $d_{y}\times d_{y}$ matrices $\{R^{d_{y}}\}_{d_{y}\in\mathbb{N}}$ and fix a constant $a>0$. If we let $\lambda_{1}\geq\cdots\geq\lambda_{d_{y}}>0$ denote the eigenvalues of $R^{d_{y}}$, arranged in decreasing order, then
$a^{d_{y}}|R^{d_{y}}|=a^{d_{y}}\prod_{i=1}^{d_{y}}\lambda_{i}.$
@@ 4.1.1 Linear Gaussian model with arbitrary $d_{x}$ and increasing $d_{y}$
If there is some constant $b>0$, independent of $d_{y}$, such that $|R^{d_{y}}|\geq a^{-d_{y}}b$, then we easily see that
$\lim_{d_{y}\to\infty}a^{d_{y}}|R^{d_{y}}|=\lim_{d_{y}\to\infty}a^{d_{y}}\prod_{i=1}^{d_{y}}\lambda_{i}\geq b>0.$ (33)
From (33), it is straightforward to provide sufficient conditions in terms of the eigenvalues of matrix $R$ that guarantee $\lim_{d_{y}\to\infty}a^{d_{y}}|R^{d_{y}}|>0$. Specifically:
  • (i)
    If $\lambda_{d_{y}}\geq\frac{1}{a}$ for all $d_{y}$, then $|R^{d_{y}}|a^{d_{y}}\geq 1$ and
    $\lim_{d_{y}\to\infty}a^{d_{y}}|R^{d_{y}}|\geq 1.$
  • (ii)
    Moreover, if there exists $\delta>0$ such that
    $\lambda_{d_{y}}\geq\frac{1+\delta}{a}\quad\text{for all }d_{y},$
    then
    $\lim_{d_{y}\to\infty}a^{d_{y}}|R^{d_{y}}|\geq\lim_{d_{y}\to\infty}(1+\delta)^{d_{y}}=\infty.$
By inspection of (32), we notice that the constant of interest is $a=2^{\frac{2}{3}}\pi$. In case (i) above, if $\lambda_{d_{y}}\geq 2^{-\frac{2}{3}}\pi^{-1}$, then
$\lim_{d_{y}\to\infty}\left(\left(2^{\frac{2}{3}}\pi\right)^{d_{y}}|R^{d_{y}}|\right)^{\frac{3}{2}}\geq 1,\;\;\implies\lim_{d_{y}\to\infty}{\mathcal{K}}_{2}<\infty,$
and, using Corollary 3.4, the approximation error can be upper bounded as
$\lim_{d_{y}\to\infty}\|\pi_{Y}(f)-\pi_{Y}^{N}(f)\|_{2}\leq\frac{B_{p}\|f\|_{\infty}}{\sqrt{N}},$
where the constant $B_{p}$ is independent of $d_{x}$, $d_{y}$ and $N$.
Furthermore, in case (ii), we obtain
$\lim_{d_{y}\to\infty}\|\pi_{Y}(f)-\pi_{Y}^{N}(f)\|_{2}\leq\lim_{d_{y}\to\infty}\frac{B_{p}\|f\|_{\infty}}{(1+\delta)^{d_{y}}\sqrt{N}}=0,$
i.e., the approximation error converges towards 0 as $d_{y}\to\infty$. This is the case for a “sufficiently regular” linear and Gaussian model, where increasing the dimension of the observations always improves the estimation.
@@
As discussed previously, the key quantity governing the dependence of the error bounds on the state dimension $d_{x}$ is the moment condition (20). However, verifying this condition directly is generally difficult in nonlinear or non-Gaussian models, since it involves high-dimensional integrals. In the following subsection, we show that, under suitable pointwise domination assumptions on the likelihood, the condition can instead be verified through simple bounds. These assumptions yield explicit estimates for the second moment of the link function required in Theorem 3.3.
@@ 4.2 Pointwise Bounded Likelihood
Consequently, if (34) and (35) hold, the second moment is finite, implying that $\ell^{\,d_{x}}\in L^{2}(\mathcal{Y};L^{2}(\pi_{0}))$. Furthermore, if there exists a uniform constant $M$ such that $M(d_{x})\leq M$ for all $d_{x}\in\mathbb{N}$, then the requirements of Corollary 3.4 are satisfied for the class of models described in Section 4.2.
The second-moment condition (20) generally involves high-dimensional integrals, which may be difficult to verify directly in practical models. In the following subsection, we bridge the gap between this abstract functional-analytic condition and concrete Bayesian modelling by deriving a tractable criterion for the integrability of the link function. Specifically, we consider likelihoods belonging to the class of elliptically symmetric families [15], which include a wide range of noise models such as Gaussian, Laplace, Student-$t$, and Cauchy distributions. For this class of models, we show that, when the observation function is bounded, the Bochner integrability condition reduces to the finiteness of a one-dimensional radial integral. This reduction enables explicit analytical constructions showing that the approximation error grows at most polynomially with the state dimension $d_{x}$. In particular, we identify scenarios in which the polynomial degree is zero, yielding error bounds that hold uniformly in $d_{x}$.
@@ 4.3 Uniformly bounded observation function
$Y=h(X)+V,$ (37)
where $h\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\to{\mathcal{Y}}$ is a (generally nonlinear) deterministic function and $V$ is an observation noise term modelling sensor and measurement errors. We assume that $V$ is independent of $X$ and admits a density on $\mathcal{Y}$}. Conditionally on $X=x$, the observation $Y$ then admits a density $g(y\mid x).$
@@
The discrepancy between an observation $y$ and its state-dependent prediction $h(x)$ is measured through the quadratic form
$(y-h(x))^{\top}R(y-h(x)),$ (39)
@@
Representation (38) separates two modelling components:
  • the geometric structure, determined by the Mahalanobis metric induced by $R$ together with the translation $h(x)$, and
@@
All the noise models obtained in Tables A and B fit into the elliptically symmetric likelihood family in (38).
@@
Consider a family of models $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$ indexed by the state dimension $d_{x}$, where each model is of the form $\mathcal{M}^{d_{x}}=\{\pi_{0},g_{y}^{d_{x}}\}$ and $g_{y}^{d_{x}}(x)$ belongs to the class of elliptically symmetric likelihoods given by (38). Hereafter, we assume that the observation map $h\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}\to{\mathcal{Y}}$ is uniformly bounded w.r.t. the Euclidean norm on $\mathcal{Y}$} by a function $M\mathrel{\mathop{\ordinarycolon}}\mathbb{N}\mapsto[0,\infty)$, of the dimension state $d_{x}$, namely
$M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h(x)\|<\infty.$ (40)
Under this assumption, it is straightforward to show (see Appendix G.1) that
$M_{R}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h(x)\|_{R}\leq\sqrt{\lambda_{1}(R)}\,M(d_{x}).$ (41)
@@
$\mathcal{S}_{R}^{d_{y}}\mathrel{\mathop{\ordinarycolon}}=|R|^{1/2}\,\mathbb{S}^{d_{y}-1}=|R|^{1/2}\,\frac{2\pi^{d_{y}/2}}{\Gamma(d_{y}/2)},$ (42)
@@
The following result reduces the high-dimensional Bochner integrability condition on the link function $\ell^{\,d_{x}}$ to the finiteness of a one-dimensional radial integral.
@@
Let the observation likelihood $g(y\mid x)$ belong to the class of elliptically symmetric families defined in (38), and assume the observation map $h\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\mapsto\mathcal{Y}$ satisfies (40). Then the second moment of the link function $\ell_{Y}^{\,d_{x}}$ is bounded by a one-dimensional radial integral, namely,
@@
A proof is provided in Appendix G.2 (see also Appendix G.1).
@@
Let the assumptions of Theorem 4.1 hold. For the observation noise models listed in Tables A and B, the integral on the right hand side of (43) is finite and can be bounded by a $d_{x}$–dependent constant $K_{2}^{d_{x}}<\infty$. Therefore
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]\leq K_{2}^{d_{x}}.$
@@
A proof is provided in Appendix G.3.
Example: Polynomial growth of $M(d_{x})$ . To illustrate the asymptotic behavior characterized in Corollary 4.2, we consider a high-dimensional Bayesian inverse problem on a state space ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$. Let $\mathbf{x}=(x_{1},\dots,x_{d_{x}})^{\top}\in{\mathcal{X}}$. We define the observation map $h\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\mapsto\mathcal{Y}$ through a system of $d_{y}$ observation channels. For each channel $i\in\{1,\dots,d_{y}\}$, the measurement is given by
$h_{i}(\mathbf{x})=\sum_{j=1}^{d_{x}}a_{ij}\sigma(x_{j}),$ (44)

### 28 May 2026 v2

@@ 4.1.1 Analysis of the Bochner constant $K_{2}^{d_{x}}$
$C_{2}=(4\uppi)^{d_{y}/2}\frac{|\Sigma_{y}^{d_{x}}|}{|R|^{1/2}}$ (31)
and then, combining (31), (29) and (24), we readily see that
$K_{2}^{d_{x}}=\mathbb{E}\!\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=\frac{|\Sigma_{y}^{d_{x}}|}{|R|},$ (32)
which is always finite (provided that $R\succ 0$, which implies $\Sigma_{y}^{d_{x}}\succ 0$ as well), ensuring that the moment condition (17) holds and so does the Bochner integrability condition (12). By (18) and (32) the Bochner constant is
$K_{2}^{d_{x}}=\frac{\left|A^{d_{x}}\Sigma_{x}^{d_{x}}(A^{d_{x}})^{\top}+R\right|}{|R|}.$ (33)
Since $R$ is positive definite, we can rewrite this expression as
$K_{2}^{d_{x}}=\left|I_{d_{y}}+R^{-1/2}A^{d_{x}}\Sigma_{x}^{d_{x}}(A^{d_{x}})^{\top}R^{-1/2}\right|.$
4.1.2 Dependence of $K_{2}^{d_{x}}$ on the dimensions $d_{x}$ and $d_{y}$
Let us define the matrix
$F^{d_{x}}=R^{-1/2}A^{d_{x}}\Sigma_{x}^{d_{x}}(A^{d_{x}})^{\top}R^{-1/2}.$
Then,
$K_{2}^{d_{x}}=|I_{d_{y}}+F^{d_{x}}|.$
Since $F^{d_{x}}$ is positive semidefinite, all its eigenvalues satisfy
$\lambda_{i}(d_{x},d_{y})\geq 0,\qquad i=1,\ldots,d_{y},$
where the eigenvalues depend on both the state dimension $d_{x}$ and the observation dimension $d_{y}$. Hence, we obtain
$K_{2}^{d_{x}}=\prod_{j=1}^{d_{y}}\left(1+\lambda_{j}(d_{x},d_{y})\right).$ (34)
Moreover, the quantity in (34) can be bounded in terms of the model parameters. Since $\Sigma_{x}^{d_{x}}$ is positive semidefinite, we may write
$F^{d_{x}}=G^{d_{x}}(G^{d_{x}})^{\top},$
where
$G^{d_{x}}=R^{-1/2}A^{d_{x}}(\Sigma_{x}^{d_{x}})^{1/2}\in\mathbb{R}^{d_{y}\times d_{x}}.$
Therefore,
$\lambda_{i}(d_{x},d_{y})=\sigma_{i}^{2}(G^{d_{x}}),\qquad i=1,\ldots,d_{y}.$
Consequently, the eigenvalues depend on the model matrices through the singular values of the noise-normalized and prior-covariance-weighted observation matrix
$R^{-1/2}A^{d_{x}}(\Sigma_{x}^{d_{x}})^{1/2}.$
We can therefore write
$K_{2}^{d_{x}}=\prod_{i=1}^{d_{y}}\left(1+\sigma_{i}^{2}(G^{d_{x}})\right).$ (35)
@@ 4.1.2 Dependence of $K_{2}^{d_{x}}$ on the dimensions $d_{x}$ and $d_{y}$
Assume that $\sigma_{i}(G^{d_{x}})\neq 0$ for all $i=1,\ldots,d_{y}$. Taking the limit as $d_{y}\to\infty$, we obtain
$\lim_{d_{y}\to\infty}K_{2}^{d_{x}}=\prod_{i=1}^{\infty}\left(1+\sigma_{i}^{2}(G^{d_{x}})\right).$
By Theorem 8.52 in [4], this infinite product converges if and only if
$\sum_{i=1}^{\infty}\sigma_{i}^{2}(G^{d_{x}})<\infty.$
Therefore, the Bochner constant is not necessarily unbounded as the observation dimension $d_{y}\to\infty$. A uniform error bound may still be obtained provided the singular values associated with the model parameters are suitably controlled.
We now return to the setting in which $d_{y}$ is fixed. Using the singular-value inequality
$\sigma_{i}(BC)\leq\sigma_{1}(B)\sigma_{i}(C),$
which holds for any pair of compatible matrices $B$ and $C$, we obtain
$\sigma_{i}^{2}(G^{d_{x}})$ $\leq\lambda_{1}(R^{-1})\sigma_{i}^{2}(A^{d_{x}})\lambda_{1}(\Sigma_{x}^{d_{x}}).$
Hence,
$\sigma_{i}^{2}(G^{d_{x}})\leq\frac{\lambda_{1}(\Sigma_{x}^{d_{x}})}{\lambda_{d_{y}}(R)}\,\sigma_{i}^{2}(A^{d_{x}}).$ (36)
Substituting (36) into (35) yields
$K_{2}^{d_{x}}\leq\prod_{i=1}^{d_{y}}\left(1+\frac{\lambda_{1}(\Sigma_{x}^{d_{x}})}{\lambda_{d_{y}}(R)}\sigma_{i}^{2}(A^{d_{x}})\right)\leq\left(1+\frac{\lambda_{1}(\Sigma_{x}^{d_{x}})}{\lambda_{d_{y}}(R)}\sigma_{1}^{2}(A^{d_{x}})\right)^{d_{y}}.$ (37)
Therefore, the dependence of the Bochner constant on the state dimension $d_{x}$ is determined by the growth of the maximal eigenvalue of $\Sigma_{x}^{d_{x}}$ and the maximal singular value of $A^{d_{x}}$.
Remark 8.
Assume that there exist polynomials $P_{n_{A}}(d_{x})$ and $P_{n_{\Sigma}}(d_{x})$, of degrees $n_{A},n_{\Sigma}\in\mathbb{N}$, such that
$\sigma_{1}(A^{d_{x}})\leq P_{n_{A}}(d_{x}),\qquad\lambda_{1}(\Sigma_{x}^{d_{x}})\leq P_{n_{\Sigma}}(d_{x}).$
Then,
$K_{2}^{d_{x}}\leq P_{n}(d_{x}),$
where $P_{n}$ is a polynomial of degree
$n=(2n_{A}+n_{\Sigma})d_{y}.$
Hence, if the maximal singular value of the observation operator and the maximal eigenvalue of the prior covariance matrix grow at most polynomially in $d_{x}$, then the Bochner constant $K_{2}^{d_{x}}$ also grows at most polynomially satisfying the assumptions in Theorem 3.3.
Moreover, if
$C_{A}\mathrel{\mathop{\ordinarycolon}}=\sup_{d_{x}\in\mathbb{N}}\sigma_{1}(A^{d_{x}})<\infty,\qquad C_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\sup_{d_{x}\in\mathbb{N}}\lambda_{1}(\Sigma_{x}^{d_{x}})<\infty,$ (38)
@@
then there exists a constant $\mathcal{K}_{2}<\infty$, independent of $d_{x}$, such that
$K_{2}^{d_{x}}\leq\mathcal{K}_{2},$
with
$\mathcal{K}_{2}\mathrel{\mathop{\ordinarycolon}}=\left(1+\frac{C_{\Sigma}\,C_{A}^{2}}{\lambda_{d_{y}}(R)}\right)^{d_{y}}.$ (39)
In particular, under these uniform boundedness assumptions, the conditions of Corollary 3.4 are satisfied. A numerical example illustrating polynomial and uniform error bounds w.r.t. the state dimension $d_{x}$ is provided in Appendix G.
@@
As discussed above, the key quantity governing the dependence of the error bounds on the state dimension $d_{x}$ is the moment condition (17). However, verifying this condition directly is generally difficult in nonlinear or non-Gaussian models, since it involves high-dimensional integrals. In the following subsection, we show that, under suitable pointwise domination assumptions on the likelihood, the condition can instead be verified through simple bounds. These assumptions yield explicit estimates for the second moment of the link function required in Theorem 3.3.
@@ 4.2 Pointwise Bounded Likelihood
Consequently, if (40) and (41) hold, the second moment is finite, implying that the link function $\ell^{\,d_{x}}\in L^{2}(\mathcal{Y};L^{2}(\pi_{0}))$. Furthermore, if there exists a uniform constant $M$ such that $M(d_{x})\leq M$ for all $d_{x}\in\mathbb{N}$, then the requirements of Corollary 3.4 are satisfied for the class of models described in Section 4.2.
The second-moment condition (17) generally involves high-dimensional integrals, which may be difficult to verify directly in practical models. In the following subsection, we bridge the gap between this abstract functional-analytic condition and concrete Bayesian modeling by deriving a tractable criterion for the integrability of the link function. Specifically, we consider likelihoods belonging to the class of elliptically symmetric families [16], which include a wide range of noise models such as Gaussian, Laplace, Student-$t$, and Cauchy distributions. For this class of models, we show that, when the observation function is bounded, the Bochner integrability condition reduces to the finiteness of a one-dimensional radial integral. This reduction enables explicit analytical constructions showing that the approximation error grows at most polynomially with the state dimension $d_{x}$. In particular, we identify scenarios in which the polynomial degree is zero, yielding error bounds that hold uniformly in $d_{x}$.
@@ 4.3 Uniformly bounded observation function
$Y=h^{d_{x}}(X)+V,$
where $h^{dx}\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\to{\mathcal{Y}}$ is a (generally nonlinear) deterministic function and $V$ is an observation noise term modeling sensor and measurement errors. We assume that $V$ is independent of $X$ and admits a density on $\mathcal{Y}$}. Conditionally on $X=x$, the observation $Y$ then admits a density $g^{d_{x}}(y\mid x).$
@@
The discrepancy between an observation $y$ and its state-dependent prediction $h^{d_{x}}(x)$ is measured through the quadratic form
$(y-h^{d_{x}}(x))^{\top}R(y-h^{d_{x}}(x)),$
@@
Representation (43) separates two modeling components:
  • the geometric structure, determined by the Mahalanobis metric induced by $R$ together with the translation $h^{d_{x}}(x)$, and
@@
All the noise models listed in Tables A and B fit into the elliptically symmetric likelihood family describe by (43).
@@
Consider a family of models $\{\mathcal{M}^{d_{x}}\}_{d_{x}\in\mathbb{N}}$ indexed by the state dimension $d_{x}$, where each model is of the form $\mathcal{M}^{d_{x}}=\{\pi_{0},g_{y}^{d_{x}}\}$ and $g_{y}^{d_{x}}(x)$ belongs to the class of elliptically symmetric likelihoods given by (43). Hereafter, we assume that the observation map $h^{d_{x}}\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}\to{\mathcal{Y}}$ is uniformly bounded w.r.t. the Euclidean norm on $\mathcal{Y}$} by a function $M\mathrel{\mathop{\ordinarycolon}}\mathbb{N}\mapsto[0,\infty)$, of the state dimension $d_{x}$, namely
$M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h^{d_{x}}(x)\|<\infty.$ (44)
Under this assumption, it is straightforward to show (see Appendix H.1) that
$M_{R}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in{\mathcal{X}}}\|h^{d_{x}}(x)\|_{R}\leq\sqrt{\lambda_{1}(R)}\,M(d_{x}).$
@@
$\mathcal{S}_{R}^{d_{y}}\mathrel{\mathop{\ordinarycolon}}=|R|^{1/2}\,\mathbb{S}^{d_{y}-1}=|R|^{1/2}\,\frac{2\pi^{d_{y}/2}}{\Gamma(d_{y}/2)},$
@@
The following result reduces the high-dimensional Bochner integrability condition for the link function $\ell^{\,d_{x}}$, given in (17), to the finiteness of a one-dimensional radial integral.
@@
Let the observation likelihood $g^{d_{x}}(y\mid x)$ belong to the class of elliptically symmetric families defined by (43). Assume that the observation map $h^{d_{x}}\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\mapsto\mathcal{Y}$ satisfies (44). Then the second moment of the link function $\ell_{Y}^{\,d_{x}}$ is bounded by a one-dimensional radial integral, namely,
@@
A proof is provided in Appendix H.2 (see also Appendix H.1).
@@
Let the assumptions of Theorem 4.1 hold. For the observation noise models listed in Tables A and B, the integral on the right hand side of (45) is finite. Therefore
$\mathbb{E}\left[\|\ell_{Y}^{\,d_{x}}\|_{L^{2}(\pi_{0})}^{2}\right]=K_{2}^{d_{x}}<\infty.$
@@
A proof is provided in Appendix H.3.
4.3.1 Example: Polynomial growth of $M(d_{x})$
To illustrate the asymptotic behavior characterized in Corollary 4.2, we consider a high-dimensional Bayesian inverse problem on a state space ${\mathcal{X}}\subseteq\mathbb{R}^{d_{x}}$. Let $\mathbf{x}=(x_{1},\dots,x_{d_{x}})^{\top}\in{\mathcal{X}}$. We define the observation map $h^{d_{x}}\mathrel{\mathop{\ordinarycolon}}{\mathcal{X}}\mapsto\mathcal{Y}$ through a system of $d_{y}$ observation channels. For each channel $i\in\{1,\dots,d_{y}\}$, the measurement is given by
$h_{i}^{d_{x}}(\mathbf{x})=\sum_{j=1}^{d_{x}}a_{ij}\sigma(x_{j}),$